OpenStudy (jjuden):

What is the length of the hypotenuse of a right triangle whose legs measure 25 inches and 60 inches? a 35 in. b 65 in. c 75 in. d 85 in.

OpenStudy (primeralph):

Use Pythagor......

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (jjuden):

thanks for helping so much (as she said sarcastically, because if she knew how to do it she would not ask for help)

OpenStudy (primeralph):

You should really calm down and learn to communicate your mind well.

OpenStudy (jjuden):

calm down i am i calm just hate people

OpenStudy (primeralph):

Well then don't come here expecting people to help.

OpenStudy (anonymous):

ok since you are finding the hypotenuse. that means we need to find C in the pathagorean theorem. it'll look like this \[60^2+25^2=c^2\]

OpenStudy (anonymous):

next u square them \[60^2+25^2=c^2 \] \[3600+625=c^2\] do you know what to do after that?

OpenStudy (anonymous):

\[4225=c^2\] get rid of the square in c by squaring 4225. That'll get you \[\sqrt{4225}\]=65